# S-EPOA: Overcoming the Indistinguishability of Segments with Skill-Driven Preference-Based Reinforcement Learning

### v2

@@ 1 Introduction
Reinforcement Learning (RL) has made significant progress across a variety of fields, including gameplay Mnih et al. (2013); Silver et al. (2016), robotics Chen et al. (2022) and autonomous systems Bellemare et al. (2020). Yet, the success of RL often relies on the careful construction of reward functions, a process that can be both labor-intensive and costly. To solve this issue, Preference-based Reinforcement Learning (PbRL) emerges as a compelling alternative Christiano et al. (2017); Lee et al. (2021b). PbRL uses human-provided preferences among various agent behaviors to serve as the reward signal, thereby eliminating the need for hand-crafted reward functions.
@@ 5.2 Results on Benchmark Tasks
Figure 3 shows the learning curves of S-EPOA and baselines on the four DMControl tasks with error rates $\epsilon\in{0.1,0.2,0.3}$. S-EPOA outperforms baselines in most environments and remains robust under non-ideal conditions, while other PbRL methods are unstable and even fail.
@@ 6 Related Work
PbRL enables humans (or supervisors in other forms, like script teachers) to guide the RL agent toward desired behaviors by providing preference on segment pairs, where the feedback efficiency is a primary concern Lee et al. (2021b); Park et al. (2022a). Prior works improve the feedback efficiency from various perspectives. Some works focus on the query selection scheme, trying to improve the information quality of queries Ibarz et al. (2018); Biyik et al. (2020); Hejna III and Sadigh (2023). Some works integrate unsupervised pretraining to avoid the waste on initial nonsense queries Lee et al. (2021b). Some works augment queries from humans to better utilize limited human feedback Park et al. (2022a); Liu et al. (2023). These methods depend on reliable feedback. However, humans could make mistakes, especially when the segment pair for comparison is slightly different, which restricts and even harms the performance in practice Lee et al. (2021a); Cheng et al. (2024).
@@ Appendix A Proof
As in Huber (2020), we approximate sigmoid function with probit function using input scaling factor $\lambda=\sqrt{\pi/8}$ Kristiadi et al. (2020), leading to an approximation of $\mathrm{Var}[\hat{P}[\sigma_{1}\succ\sigma_{2}]]$:
@@ B.1 Tasks
    Door_open: An agent controls a robotic arm to open a door with random position.
@@
    Button_press: An agent controls a robotic arm to press a button with random position.
@@
    Window_open: An agent controls a robotic arm to open a window with random position.
@@ B.2 Implementation Details
The APS, DIAYN and CIC algorithms used in S-EPOA are implemented based on the official repository of URLB Laskin et al. (2021). We changed the DDPG backbone of URLB to SAC, for PEBBLE, SURF, and RUNE are all based on SAC. Also, for APS, to encourage exploration in the pretraining stage, we add a hyperparameter $\beta>1$ for the entropy term in intrinsic reward (Eq. 13):
@@
For the implementation of SAC, PEBBLE, SURF and RUNE, we refer to their corresponding official repositories, and re-implement them within the URLB framework. For RIME, we use its official implementation. SAC serves as a performance upper bound because it uses a ground-truth reward function which is unavailable in PbRL settings for training. The detailed hyperparameters of SAC are shown in Table 3. PEBBLE’s settings remain consistent with its original implementation, and the specifics are detailed in Table 4. For SURF, RUNE, RIME and S-EPOA, most hyperparameters are the same as those of PEBBLE, and other hyperparameters are detailed in Table 5, 6, 7 and 8, respectively. The total amount of feedback and feedback amount per session are detailed in Table 9.
For the component analysis of ablation studies in Section. 5.3, we remove skill-based unsupervised pretraining from S-EPOA (green and blue curves) by setting the pretraining step = 0 and randomly initialize the APS policy. As for the skill-based query selection, we replace it with the disagreement scheme in PEBBLE and SURF (orange and blue curves).
We have anonymously opened sourced the implementation of S-EPOA  [^1] Anonymous repository URL: https://anonymous.4open.science/r/SEPOA_PbRL.

### v3

@@ 1 Introduction
Reinforcement Learning (RL) has made significant progress across a variety of fields, including gameplay Mnih et al. (2013); Silver et al. (2016), robotics Chen et al. (2022) and autonomous systems Bellemare et al. (2020); Mu et al. (2024); Luan et al. (2025). Yet, the success of RL often relies on the careful construction of reward functions, a process that can be both labor-intensive and costly. To solve this issue, Preference-based Reinforcement Learning (PbRL) emerges as a compelling alternative Christiano et al. (2017); Lee et al. (2021b). PbRL uses human-provided preferences among various agent behaviors to serve as the reward signal, thereby eliminating the need for hand-crafted reward functions.
@@ 5.2 Results on Benchmark Tasks
Figure 3 shows the learning curves of S-EPOA and baselines on the four DMControl tasks with error rates $\epsilon\in\{0.1,0.2,0.3\}$. S-EPOA outperforms baselines in most environments and remains robust under non-ideal conditions, while other PbRL methods are unstable and even fail.
@@ 6 Related Work
PbRL enables humans (or supervisors in other forms, like script teachers) to guide the RL agent toward desired behaviors by providing preferences on segment pairs, where the feedback efficiency is a primary concern Lee et al. (2021b); Park et al. (2022a). Prior works improve the feedback efficiency from various perspectives. Some works focus on the query selection scheme, trying to improve the information quality of queries Ibarz et al. (2018); Biyik et al. (2020); Hejna III and Sadigh (2023). Some works integrate unsupervised pretraining to avoid the waste on initial nonsense queries Lee et al. (2021b). Some works augment queries from humans to better utilize limited human feedback Park et al. (2022a); Liu et al. (2023). These methods depend on reliable feedback. However, humans could make mistakes, especially when the segment pair for comparison is slightly different, which restricts and even harms the performance in practice Lee et al. (2021a); Cheng et al. (2024).
@@ 7 Conclusion
Acknowledgments
This work is supported by the NSFC (No. 62125304, 62192751, and 62073182), the Beijing Natural Science Foundation (L233005), BNRist project (BNR2024TD03003), and the 111 International Collaboration Project (B25027).
@@ Appendix A Proof
As in Huber (2020), we approximate the sigmoid function with the probit function using the input scaling factor $\lambda=\sqrt{\pi/8}$ Kristiadi et al. (2020), leading to an approximation of $\mathrm{Var}[\hat{P}[\sigma_{1}\succ\sigma_{2}]]$:
@@ B.1 Tasks
    Door_open: An agent controls a robotic arm to open a door with a random position.
@@
    Button_press: An agent controls a robotic arm to press a button at a random position.
@@
    Window_open: An agent controls a robotic arm to open a window with a random position.
@@ B.2 Implementation Details
The APS, DIAYN and CIC algorithms used in S-EPOA are implemented based on the official repository of URLB Laskin et al. (2021). We changed the DDPG backbone of URLB to SAC, as PEBBLE, SURF, and RUNE are all based on SAC. Also, for APS, to encourage exploration in the pretraining stage, we add a hyperparameter $\beta>1$ for the entropy term in intrinsic reward (Eq. 13):
@@
For the implementation of SAC, PEBBLE, SURF, and RUNE, we refer to their corresponding official repositories and re-implement them within the URLB framework. For RIME, we use its official implementation. SAC serves as a performance upper bound because it uses a ground-truth reward function, which is unavailable in PbRL settings for training. The detailed hyperparameters of SAC are shown in Table 3. PEBBLE’s settings remain consistent with its original implementation, and the specifics are detailed in Table 4. For SURF, RUNE, RIME and S-EPOA, most hyperparameters are the same as those of PEBBLE, and other hyperparameters are detailed in Table 5, 6, 7 and 8, respectively. The total amount of feedback and feedback amount per session are detailed in Table 9.
For the component analysis of ablation studies in Section. 5.3, we remove skill-based unsupervised pretraining from S-EPOA (green and blue curves) by setting the pretraining step = 0 and randomly initializing the APS policy. As for the skill-based query selection, we replace it with the disagreement scheme in PEBBLE and SURF (orange and blue curves).
Code of S-EPOA is available at https://github.com/MoonOutCloudBack/SEPOA_PbRL.